OpenStudy (anonymous):

Input the equation of the given line in slope-intercept form. The line through (1, 0) and parallel to x - y = 7.

OpenStudy (sasogeek):

if parallel, then slopes are equal. slope of given equation x-y=7 is 1 taking any point P(x,y) on the line passing through (1,0) => slope = y-0/x-1 =1 y-0=(x-1)1 y=x-1